# Task-Oriented Feature Compression for Multimodal Understanding via Device-Edge Co-Inference

### v1

@@ Abstract
With the rapid development of large multimodal models (LMMs), multimodal understanding applications are emerging. As most LMM inference requests originate from edge devices with limited computational capabilities, the predominant inference pipeline involves directly forwarding the input data to an edge server which handles all computations. However, this approach introduces high transmission latency due to limited uplink bandwidth of edge devices and significant computation latency caused by the prohibitive number of visual tokens, thus hindering delay-sensitive tasks and degrading user experience. To address this challenge, we propose a task-oriented feature compression (TOFC) method for multimodal understanding in a device-edge co-inference framework, where visual features are merged by clustering and encoded by a learnable and selective entropy model before feature projection. Specifically, we employ density peaks clustering based on $K$ nearest neighbors to reduce the number of visual features, thereby minimizing both data transmission and computational complexity. Subsequently, a learnable entropy model with hyperprior is utilized to encode and decode merged features, further reducing transmission overhead. To enhance compression efficiency, multiple entropy models are adaptively selected based on the characteristics of the visual features, enabling a more accurate estimation of the probability distribution. Comprehensive experiments on seven visual question answering benchmarks validate the effectiveness of the proposed TOFC method. Results show that TOFC achieves up to 60% reduction in data transmission overhead and 50% reduction in system latency while maintaining identical task performance, compared with traditional image compression methods.
Index Terms:
Distributed inference, large multimodal models, edge artificial intelligence, task-oriented communication.
@@ I Introduction
Large multimodal models (LMMs) have recently emerged as a powerful tool by integrating multimodal understanding capabilities with versatile large language models (LLMs) [1, 2]. LMMs have the potential to be applied to a variety of domains, including embodied artificial intelligence [3], smart manufacturing, autonomous driving [4] and personal assistants [5]. The architecture of mainstream LMMs for vision input comprises three components: a vision encoder, a feature projector and an LLM [6, 7, 8]. The vision encoder is typically based on contrastive language–image pretraining (CLIP) [9], which transforms image patches into visual features with lingual knowledge, benefiting from large-scale pretraining on image-text pairs. Subsequently, the feature projector maps the extracted features into visual tokens, thus bridging the gap between CLIP and the LLM. Finally, the LLM processes both visual and textual tokens and generates response in natural language. This architecture leverages the prior knowledge of both CLIP and LLM, accumulated from extensive pretraining, thereby reducing the training cost [6]. However, a fundamental limitation of CLIP-based vision encoders [9, 10] is that they represent a small image patch by a large number of features, leading to a prohibitive number of visual tokens. As the computational complexity of attention mechanisms, which are widely adopted in LLMs, grows quadratically with the length of input sequence, the auto-regressive inference of LLMs incurs significant latency. This challenge is particularly critical for applications that require low-latency responses and real-time interactions, such as analyzing surrounding environment for autonomous robots [3]. Moreover, most inference requests originate from edge devices with limited computational capabilities. Consequently, the predominant LMM inference pipeline involves directly transmitting the input data to a proximate edge server, which handles all computations. However, in scenarios where only the wireless network connection is available for the user device, the communication channel is highly dynamic and the uplink bandwidth is often stringent. The transmission of visual data over such channel incurs substantial latency, which hinders the execution of delay-sensitive tasks [11]. Therefore, the data communication overhead and high inference latency constitute the main design challenge of LMM inference.
@@
To reduce bandwidth consumption and response time in LMM inference, a device-edge co-inference framework [12] stands out as a promising solution. One co-inference paradigm under heated discussion is task-oriented communication [11, 13, 14, 15], which designs the transmission strategies based on the characteristics of downstream tasks rather than recovering the original data at the edge server, thus representing a paradigm shift from traditional data-oriented approaches. In task-oriented co-inference schemes, the user device extracts task-relevant information and transmits only the essential features to the edge server, which handles the extensive computation of model inference. While the device-edge co-inference framework with task-oriented communication has been explored for several vision tasks [16, 17, 18], these methods are not directly applicable to LMM inference. The primary challenge lies in the fact that LMMs rely on the CLIP model, which is pretrained on large-scale datasets, to extract visual features necessary for diverse downstream tasks. Given this dependency, training an alternative vision encoder with constrained communication overhead is impractical.
@@
Another line of research aims to accelerate LMM inference by reducing the number of visual tokens, thereby minimizing the end-to-end system latency. These methods either prune unimportant visual tokens or merge multiple visual tokens into a single representation at different stages of LMM inference, including vision encoding [26, 27], feature projection [28, 29, 30] and LLM generation [31]. However, these methods assume that the entire inference process is executed on a single device and do not consider the constraints on data transmission in a device-edge co-inference framework. Given that latent features remain high-dimensional throughout LMM inference, these methods are not suitable for co-inference schemes without the incorporation of other feature compression techniques.
@@ I-A Contributions
  • The effectiveness of the proposed TOFC method is validated through comprehensive experiments on seven visual question answering (VQA) benchmarks. Results show that our TOFC method achieves substantial reductions in both data transmission overhead and system latency. Notably, on the RealWorldQA benchmark, our TOFC method reduces data transmission overhead by up to 60% and system latency by 50% while maintaining identical task performance compared with traditional JPEG compression. In addition, case studies are conducted to intuitively demonstrate the advantages of our TOFC method, and an ablation study validates the effectiveness of the proposed FMM and entropy model.
@@ II-A Task-Oriented Communication
The increasing demand for multimodal services has stimulated intensive interactions between user devices and the edge server, resulting in a surge in multimodal data transmission and posing severe challenges to existing communication systems. Conventional data-oriented transmission paradigm focuses on reliable transmission of every single bit of multimodal input, while overlooking the inherent redundancy in the data. To address this limitation, task-oriented communication [11, 13, 14, 15] proposes to extract the task-related information, thereby reducing data transmission overhead while maintaining satisfactory performance on downstream tasks. This objective is achieved by adopting the information bottleneck (IB) principle [32, 33, 34], which maximizes task performance while minimizing communication overhead. Several recent works have leveraged task-oriented communication to reduce transmission overhead by over an order of magnitude. For instance, the authors in [35] establish a temporal entropy model to minimize the redundancy in sequential video frames while achieving exceptional video analysis performance. Follow-up studies investigate on extending task-oriented principle to other scenarios of mobile communications, including multi-device cooperative inference [16], image processing with digital modulation [17], domain generalization [36], and cooperative perception in autonomous driving [18]. The above studies primarily investigate conventional vision tasks such as object detection and classification, while no existing work has addressed the challenge of prohibitive transmission overhead in LMM inference. Given the surging demands for LMM inference, we adopt the task-oriented communication principle in multimodal understanding tasks. As LMMs rely on high-dimensional visual inputs and large neural networks to achieve state-of-the-art performance across diverse and complex multimodal tasks, applying the task-oriented principle to LMM inference constitutes a unique challenge.
@@ II-B Neural Data Compression
Various data compression methods have been proposed to minimize storage and transmission overhead, especially for high-dimensional visual data. Traditional visual data compression methods, such as JPEG [19, 20], BPG [37], and WebP [21], adopt linear transforms [38] to obtain compact representations in the frequency domain, followed by entropy coding to further remove redundancy. In contrast, neural compression methods [22, 23, 24] leverage neural networks to extract latent features and optimize the trade-off between data rate and reconstruction quality. Benefiting from the versatile nonlinear transforms [39] and end-to-end optimization, these learning-based methods outperform traditional compression methods. Subsequent studies [40, 41, 42, 25] equip the learnable entropy model with autoregressive structures to extract correlated contexts in both spatial and channel dimensions. These contexts exploit the intrinsic correlations in latent features to obtain a more accurate estimation of the statistics of latent features, thereby enhancing the rate-distortion (RD) performance. However, previous data compression methods assess reconstruction quality by traditional metrics such as peak signal-to-noise ratio (PSNR), focusing on high-fidelity image reconstruction in the pixel space. This design encourages the model to capture general features in the visual data instead of extracting specific task-relevant information, resulting in information redundancy and excessive communication overhead. To improve transmission efficiency, this work considers feature compression based on task-oriented communication principles, which eliminates the task-irrelevant redundancy in the visual data by directly optimizing the trade-off between transmission overhead and performance on downstream tasks.
@@ III System Model
As illustrated in Fig. 2, we propose a device-edge co-inference system for LMMs, where the user device processes and encodes the input image before transmitting the encoded visual features, along with the textual instruction, to an edge server for LMM inference. Specifically, the user device takes input of two modalities: a textual instruction that describes the user’s request in natural language and a visual input that contains semantic information necessary for downstream tasks. The textual instruction is directly transmitted to the edge server due to its small data size. On the other hand, the input image accounts for the majority of the transmission overhead. Each image is represented as a tensor of pixel-wise RGB intensities $\boldsymbol{v}_{\rm i}\in\mathbb{R}^{h\times w\times 3}$, where $h$ and $w$ represent the height and width of the image, respectively. To preserve the fine-grained details in the input image, it is padded and partitioned into $n_{\rm h}\times n_{\rm w}$ patches, each with a resolution of $p\times p$ that satisfies the requirement of the vision encoder [7]. Additionally, a thumbnail of the input image is created by padding the original image to a square shape and resizing it to the size of one patch as $p\times p$, which captures the overall structure of the image [7]. The visual signal consists of both the detailed patches and the overall thumbnail, written as $\boldsymbol{v}_{\rm p}\in\mathbb{R}^{n_{\rm p}\times p\times p\times 3}$, where $n_{\rm p}=n_{\rm h}\times n_{\rm w}+1$ denotes the total number of patches.
@@
where $f_{\rm vision}(\cdot)$ denotes the vision encoder and $\boldsymbol{x}\in\mathbb{R}^{n_{\rm p}\times n_{\rm v}\times d_{\rm v}}$ represents visual features, with $n_{\rm v}$ and $d_{\rm v}$ denoting the number of visual features and the dimension of each visual feature, respectively. To reduce data size and accelerate inference, we propose a feature merging module that transforms the visual features $\boldsymbol{x}$ into merged features $\boldsymbol{y}\in\mathbb{R}^{n_{\rm p}\times n_{\rm c}\times d_{\rm v}}$, where $n_{\rm c}$ is the number of merged features. These merged features are encoded into a bitstream using a learnable entropy encoder [23] and then transmitted through a wireless channel to the edge server.
At the edge server, the received textual instruction is tokenized using the codebook of the pre-trained LLM. Meanwhile, the received bitstream is decoded by an entropy decoder to recover the merged features $\hat{\boldsymbol{y}}$. Subsequently, a two-layer multilayer perceptron (MLP) projects the merged features into the word embedding space, thus aligning the visual input with the textual tokens [7]. The resulting visual tokens are concatenated with the textual tokens, constituting the input of the LLM. The tokens generated by the LLM, denoted by $\boldsymbol{h}_{\rm a}$, are transformed into plain text, and this textual response is transmitted back to the user device for downstream tasks.
@@ IV-B Learnable Entropy Model-Based Codec
[…]
@@
As the dimension of each visual feature increases significantly when transformed to the word embedding space, we employ entropy coding and transmit the bitstream before feature projection. Effective and compact entropy coding relies on accurate probability distribution of the discrete variable [39]. However, it is practically infeasible to directly estimate arbitrary probability distribution. As a solution, we model the merged features $\boldsymbol{y}$ by Laplacian distribution, where the mean and scale are derived from the hyperprior $\boldsymbol{z}$. The probability distribution of the hyperprior $\boldsymbol{z}$ is estimated by a learnable factorized density model [23]. Both $\boldsymbol{y}$ and $\boldsymbol{z}$ are quantized and encoded by an entropy coder based on the estimated probability distribution, thereby obtaining a more compact bitstream representation while maintaining acceptable distortion for downstream tasks.
As illustrated in Fig. 3, a hyperprior analysis network $h_{\rm a}$ is employed to extract the hyperprior $\boldsymbol{z}$ from the merged features $\boldsymbol{y}$. Since the entropy coder operates on integers, both $\boldsymbol{y}$ and $\boldsymbol{z}$ are quantized before entropy coding [39]. The mean and scale of the merged features are estimated from the quantized hyperprior $\bar{\boldsymbol{z}}$ using a hyperprior synthesis network $h_{\rm s}$ as:
@@
where $\boldsymbol{\mu}$ and $\boldsymbol{b}$ denote the estimated mean and scale of the merged features $\boldsymbol{y}$, respectively. The quantization of merged features introduces an additive uniform noise [22] as $\bar{\boldsymbol{y}}=\boldsymbol{y}+\boldsymbol{o}$, where $\bar{\boldsymbol{y}}$ denotes the quantized merged features, $\boldsymbol{o}\sim\mathcal{U}(-0.5,0.5)$ is the additive noise, and $\mathcal{U}(a,b)$ represents a uniform distribution from $a$ to $b$. As illustrated in Fig. 4, the empirical measurement of the probability density function (PDF) of the visual features resembles the Laplacian distribution. Consequently, the probability distribution of $\bar{\boldsymbol{y}}$ is modeled as Laplacian convolved with a uniform distribution, expressed as:
@@
where $\mathcal{L}(\boldsymbol{\mu},\boldsymbol{b})$ represents a Laplacian distribution with a mean vector of $\boldsymbol{\mu}$ and a scale vector of $\boldsymbol{b}$, and * denotes convolution. Similarly, the probability distribution of quantized hyperprior $\bar{\boldsymbol{z}}$ is written as:
@@
where $p(\boldsymbol{z})$ is the estimated probability distribution of $\boldsymbol{z}$ [23]. To allow back-propagation of gradients through the quantization operation, we employ the straight-through estimator (STE) during training [41], expressed as:
@@
where $\boldsymbol{h}_{{\rm a},t}$ denotes the $t$-th token predicted by the LLM, and $T$ is the length of the ground-truth response. The rate loss is calculated as the average combined entropy of quantized merged features and the hyperprior, which is estimated from probability distribution as [23]:
$R=-\frac{1}{n_{\rm p}n_{\rm c}d_{\rm v}}(\log_{2}p(\bar{\boldsymbol{y}}|\bar{\boldsymbol{z}})+\log_{2}p(\bar{\boldsymbol{z}})).$ (11)

### v3

@@ Abstract
With the rapid development of large multimodal models (LMMs), multimodal understanding applications are emerging. As most LMM inference requests originate from edge devices with limited computational capabilities, the predominant inference pipeline involves directly forwarding the input data to an edge server which handles all computations. However, this approach introduces high transmission latency due to limited uplink bandwidth of edge devices and significant computation latency caused by the prohibitive number of visual tokens, thus hindering delay-sensitive tasks and degrading user experience. To address this challenge, we propose a task-oriented feature compression (TOFC) method for multimodal understanding in a device-edge co-inference framework, where visual features are merged by clustering and encoded by a learnable and selective entropy model before feature projection. Specifically, we employ density peaks clustering based on $K$ nearest neighbors to reduce the number of visual features, thereby minimizing both data transmission and computational complexity. Subsequently, a learnable entropy model with hyperprior is utilized to encode and decode merged features, further reducing transmission overhead. To enhance compression efficiency, multiple entropy models are adaptively selected based on the characteristics of the visual features, enabling a more accurate estimation of the probability distribution. Comprehensive experiments on seven visual question answering benchmarks validate the effectiveness of the proposed TOFC method. Results show that TOFC achieves up to 52% reduction in data transmission overhead and 63% reduction in system latency while maintaining identical task performance, compared with neural compression ELIC.
@@ I Introduction
Large multimodal models (LMMs) have recently emerged as a powerful tool by integrating multimodal understanding capabilities with versatile large language models (LLMs) [1, 2]. LMMs have the potential to be applied to a variety of domains, including embodied artificial intelligence [3], smart manufacturing, autonomous driving [4], and personal assistants [5]. The architecture of mainstream LMMs for vision input comprises three components: a vision encoder, a feature projector, and an LLM [6, 7, 8]. The vision encoder is typically based on contrastive language–image pretraining (CLIP) [9], which transforms image patches into visual features with lingual knowledge, benefiting from large-scale pretraining on image-text pairs. Subsequently, the feature projector maps the extracted features into visual tokens, thus bridging the gap between CLIP and the LLM. Finally, the LLM processes both visual and textual tokens and generates a response in natural language. This architecture leverages the prior knowledge of both CLIP and LLM, accumulated from extensive pretraining, thereby reducing the training cost [6]. However, a fundamental limitation of CLIP-based vision encoders [9, 10] is that they represent a small image patch by a large number of features, leading to a prohibitive number of visual tokens. As the computational complexity of attention mechanisms, which are widely adopted in LLMs, grows quadratically with the length of input sequence, the auto-regressive inference of LLMs incurs significant latency. This challenge is particularly critical for applications that require low-latency responses and real-time interactions, such as analyzing the surrounding environment for autonomous robots [3]. Moreover, most inference requests originate from edge devices with limited computational capabilities. Consequently, the predominant LMM inference pipeline involves directly transmitting the input data to a proximate edge server, which handles all computations. However, in scenarios where only the wireless network connection is available for the user device, the communication channel is highly dynamic and the uplink bandwidth is often stringent. The transmission of visual data over such channel incurs substantial latency, which hinders the execution of delay-sensitive tasks [11]. Therefore, the data communication overhead and high inference latency constitute the main design challenge of LMM inference.
@@
To reduce bandwidth consumption and response time in LMM inference, a device-edge co-inference framework [12, 13] stands out as a promising solution. One co-inference paradigm under heated discussion is task-oriented communication [11, 14, 15, 16], which designs the transmission strategies based on the characteristics of downstream tasks rather than recovering the original data at the edge server, thus representing a paradigm shift from traditional data-oriented approaches. In task-oriented co-inference schemes, the user device extracts task-relevant information and transmits only the essential features to the edge server, which handles the extensive computation of model inference. While the device-edge co-inference framework with task-oriented communication has been explored for several vision tasks [17, 18, 19], these methods are not directly applicable to LMM inference. The primary challenge lies in the fact that LMMs rely on the CLIP model, which is pretrained on large-scale datasets, to extract visual features necessary for diverse downstream tasks. Given this dependency, it is impractical to train an alternative vision encoder with more compact visual features from scratch to accommodate the constrained uplink bandwidth.
@@
Another line of research aims to accelerate LMM inference by reducing the number of visual tokens, thereby minimizing the end-to-end system latency. These methods either prune unimportant visual tokens or merge multiple visual tokens into a single feature at different stages of LMM inference, including vision encoding [27, 28], feature projection [29, 30, 31] and LLM generation [32]. Given that the dimension of visual features produced by the vision encoder significantly grows after being projected to the LLM latent space [33, 7], the transmission of visual features should be conducted before MLP projection to minimize data size under identical token count. As the majority of token reduction methods are performed after MLP projection, these methods [29, 30, 31, 32] do not alleviate the data transmission overhead in a device-edge co-inference framework, which constitutes a key bottleneck in our study. Other works [27, 28] still assume that the entire inference process is executed on a single device and do not encode the visual features efficiently to meet the data transmission constraints.
@@ I-A Contributions
  • The effectiveness of the proposed TOFC method is validated through comprehensive experiments on seven visual question answering (VQA) benchmarks. Results show that our TOFC method achieves substantial reductions in both data transmission overhead and system latency. Notably, on the RealWorldQA benchmark, our TOFC method reduces data transmission overhead by up to 52% and system latency by 61% while maintaining identical task performance compared with neural compression ELIC [26]. In addition, case studies are conducted to intuitively demonstrate the advantages of our TOFC method, and an ablation study validates the effectiveness of the proposed FMM and entropy model.
@@ II-A Task-Oriented Communication
The increasing demand for multimodal services has stimulated intensive interactions between user devices and the edge server, resulting in a surge in multimodal data transmission and posing severe challenges to existing communication systems. The conventional data-oriented transmission paradigm focuses on the reliable transmission of every single bit of multimodal input, while overlooking the inherent redundancy in the data. To address this limitation, task-oriented communication [11, 14, 15, 16] proposes to extract the task-related information, thereby reducing data transmission overhead while maintaining satisfactory performance on downstream tasks. This objective is achieved by adopting the information bottleneck (IB) principle [34, 35, 36], which maximizes task performance while minimizing communication overhead. Several recent works have leveraged task-oriented communication to reduce transmission overhead by over an order of magnitude. For instance, the authors in [37] establish a temporal entropy model to minimize the redundancy in sequential video frames while achieving exceptional video analysis performance. Follow-up studies investigate extending task-oriented principle to other scenarios of mobile communications, including multi-device cooperative inference [17], image processing with digital modulation [18], domain generalization [38], and cooperative perception in autonomous driving [19]. The above studies primarily investigate conventional vision tasks such as object detection and classification, while no existing work has addressed the challenge of prohibitive transmission overhead in LMM inference. Given the surging demands for LMM inference, we adopt the task-oriented communication principle in multimodal understanding tasks. As LMMs rely on high-dimensional visual inputs and large neural networks to achieve state-of-the-art performance across diverse and complex multimodal tasks, applying the task-oriented principle to LMM inference constitutes a unique challenge.
@@ II-B Neural Data Compression
Various data compression methods have been proposed to minimize storage and transmission overhead, especially for high-dimensional visual data. Traditional visual data compression methods, such as JPEG [20, 21], BPG [39], and WebP [22], adopt linear transforms [40] to obtain compact representations in the frequency domain, followed by entropy coding to further remove redundancy. In contrast, neural compression methods [23, 24, 25] leverage neural networks to extract latent features and optimize the trade-off between data rate and reconstruction quality. Benefiting from the versatile nonlinear transforms [41] and end-to-end optimization, these learning-based methods outperform traditional compression methods. Subsequent studies [42, 43, 44, 26, 45] equip the learnable entropy model with autoregressive structures to extract correlated contexts in both spatial and channel dimensions. These contexts exploit the intrinsic correlations in latent features to obtain a more accurate estimation of the statistics of latent features, thereby enhancing the rate-distortion (RD) performance. However, previous data compression methods assess reconstruction quality by traditional metrics such as peak signal-to-noise ratio (PSNR), focusing on high-fidelity image reconstruction in the pixel space. This design encourages the model to capture general features in the visual data instead of extracting specific task-relevant information, resulting in information redundancy and excessive communication overhead. To improve transmission efficiency, this work considers latent feature compression based on task-oriented communication principles, which eliminates the task-irrelevant redundancy in the visual features by directly optimizing the trade-off between transmission overhead and performance on downstream tasks.
@@ III System Model
As illustrated in Fig. 2, we propose a device-edge co-inference system for LMMs, where the user device processes and encodes the input image before transmitting the encoded visual features, along with the textual instruction, to an edge server for LMM inference. Specifically, the user device takes input of two modalities: a textual instruction that describes the user’s request in natural language and a visual input that contains semantic information necessary for downstream tasks. The textual instruction is directly transmitted to the edge server due to its small data size. On the other hand, the input image accounts for the majority of the transmission overhead. Each image is represented as a tensor of pixel-wise RGB intensities $\boldsymbol{v}_{\rm i}\in\mathbb{R}^{h\times w\times 3}$, where $h$ and $w$ represent the height and width of the image, respectively. To preserve the fine-grained details in the input image, it is enlarged, padded and partitioned into $n_{\rm h}\times n_{\rm w}$ patches [7], each with a resolution of $p\times p$ that matches the input size of the vision encoder [10]. Additionally, a thumbnail of the input image is created by padding the original image to a square shape and resizing it to the size of one patch as $p\times p$, which captures the overall structure of the image [7]. The visual signal consists of both the detailed patches and the overall thumbnail, written as $\boldsymbol{v}_{\rm p}\in\mathbb{R}^{n_{\rm p}\times p\times p\times 3}$, where $n_{\rm p}=n_{\rm h}\times n_{\rm w}+1$ denotes the total number of patches.
@@
where $f_{\rm vision}(\cdot)$ denotes the vision encoder and $\boldsymbol{x}\in\mathbb{R}^{n_{\rm p}\times n_{\rm v}\times d_{\rm v}}$ represents visual features, with $n_{\rm v}$ and $d_{\rm v}$ denoting the number of visual features and the dimension of each visual feature, respectively. To reduce data size and accelerate inference, we propose a feature merging module that transforms the visual features $\boldsymbol{x}$ into merged features $\boldsymbol{y}\in\mathbb{R}^{n_{\rm p}\times n_{\rm c}\times d_{\rm v}}$, where $n_{\rm c}$ is the number of merged features. These merged features are encoded into a bitstream using a learnable entropy encoder [24] and then transmitted through a wireless channel to the edge server. Following [37, 17], we consider the uplink bandwidth constraints on edge devices instead of explicitly modeling the physical channel, as channel coding and signal modulation are beyond the scope of this paper.
At the edge server, the received textual instruction is tokenized using the codebook of the pre-trained LLM. Meanwhile, the received bitstream is decoded by an entropy decoder to recover the merged features $\hat{\boldsymbol{y}}$. Subsequently, a two-layer multilayer perceptron (MLP) projects the merged features into the word embedding space, thus aligning the visual input with the textual tokens [7]. The resulting visual tokens are concatenated with the textual tokens, constituting the input of the LLM. The tokens generated by the LLM, denoted by $\boldsymbol{x}_{\rm a}$, are transformed into plain text, and this textual response is transmitted back to the user device for downstream tasks.
@@ IV-B Learnable Entropy Model-Based Codec
Figure 3: Network architecture of the learnable entropy model. The hyperprior analysis network $h_{\rm{a}}$ extracts the hyperprior $\boldsymbol{z}$ from the merged features $\boldsymbol{y}$. The hyperprior synthesis network $h_{\rm s}$ estimates the mean and scale of the merged features from the quantized hyperprior $\bar{\boldsymbol{z}}$. FC represents a fully connected layer, where the two parameters in parentheses denote input and output dimensions, respectively. AE and AD represent arithmetic encoder and decoder, respectively. $\lfloor\cdot\rceil$ denotes rounding to the nearest integer.
@@
As the dimension of each visual feature increases significantly when transformed to the word embedding space, we employ entropy coding and transmit the bitstream before feature projection. Effective and compact entropy coding relies on the accurate probability distribution of the discrete variable [41]. However, it is practically infeasible to directly estimate arbitrary probability distribution. As a solution, we model the merged features $\boldsymbol{y}$ by Laplacian distribution, where the mean and scale are derived from the hyperprior $\boldsymbol{z}$. The probability distribution of the hyperprior $\boldsymbol{z}$ is estimated by a learnable factorized entropy model [24]. Both $\boldsymbol{y}$ and $\boldsymbol{z}$ are quantized and encoded by an entropy coder based on the estimated probability distribution, thereby obtaining a more compact bitstream representation while maintaining acceptable distortion for downstream tasks.
As illustrated in Fig. 3, a hyperprior analysis network $h_{\rm a}$ is employed to extract the hyperprior $\boldsymbol{z}\in\mathbb{R}^{n_{\rm p}\times n_{\rm c}\times\frac{d_{\rm v}}{16}}$ from the merged features $\boldsymbol{y}$. Since the entropy coder operates on integers, both $\boldsymbol{y}$ and $\boldsymbol{z}$ are quantized before entropy coding [41]. The mean and scale of the merged features are estimated from the quantized hyperprior $\bar{\boldsymbol{z}}$ using a hyperprior synthesis network $h_{\rm s}$ as:
@@
where $\boldsymbol{\mu}$ and $\boldsymbol{b}$ denote the estimated mean and scale of the merged features $\boldsymbol{y}$, respectively. The quantization of merged features introduces an additive uniform noise [23] as $\bar{\boldsymbol{y}}=\boldsymbol{y}+\boldsymbol{o}$, where $\bar{\boldsymbol{y}}$ denotes the quantized merged features, $\boldsymbol{o}\sim\mathcal{U}(-0.5,0.5)$ is the additive noise, and $\mathcal{U}(a,b)$ represents a uniform distribution from $a$ to $b$. As illustrated in Fig. 4, the empirical measurement of the probability density function (PDF) of the visual features resembles the Laplacian distribution. Consequently, the probability mass function (PMF) of $\bar{\boldsymbol{y}}$ is modeled as Laplacian convolved with a uniform distribution, expressed as:
@@
where $\mathcal{L}(\boldsymbol{\mu},\boldsymbol{b})$ represents a Laplacian distribution with a mean vector of $\boldsymbol{\mu}$ and a scale vector of $\boldsymbol{b}$, and * denotes convolution. Similarly, the PMF of quantized hyperprior $\bar{\boldsymbol{z}}$ is written as:
@@
where $p(\boldsymbol{z})$ is the estimated probability distribution of $\boldsymbol{z}$ [24]. Similar to prior works on vector quantized-variational autoencoder (VQ-VAE) [52] and neural image compression [43], we employ the straight-through estimator (STE) during training to allow the back-propagation of gradients through the non-differentiable quantization operation, expressed as:
@@
where $\boldsymbol{x}_{{\rm a},t}$ denotes the $t$-th token predicted by the LLM, and $T$ is the length of the ground-truth response. The rate loss should be the joint entropy of quantized merged features and the hyperprior as
$H(\bar{\boldsymbol{y}},\bar{\boldsymbol{z}})$ $=H(\bar{\boldsymbol{y}}|\bar{\boldsymbol{z}})+H(\bar{\boldsymbol{z}})$
$=\mathbb{E}_{\boldsymbol{v}_{\rm i}\sim p_{\boldsymbol{v}_{\rm i}}}[-\log p_{\bar{\boldsymbol{y}}|\bar{\boldsymbol{z}}}(\bar{\boldsymbol{y}}|\bar{\boldsymbol{z}},\boldsymbol{v}_{\rm i})-\log p_{\bar{\boldsymbol{z}}}(\bar{\boldsymbol{z}}|\boldsymbol{v}_{\rm i})],$ (11)
where $H(\cdot)$ denotes the Shannon entropy, $\mathbb{E}[\cdot]$ is the expectation operator, and $p_{\boldsymbol{v}_{\rm i}}$ represents the distribution of the input image. However, the true posterior of the visual features given the input image is assumed intractable [24], and the joint entropy can not be directly derived from (11). To address this issue, a tractable upper bound of the joint entropy is formulated based on the conditional entropy model $p(\bar{\boldsymbol{y}}|\bar{\boldsymbol{z}})$ and the factorized entropy model $p(\bar{\boldsymbol{z}})$ [37], expressed as:
$H(\bar{\boldsymbol{y}},\bar{\boldsymbol{z}})\leq\mathbb{E}_{\bar{\boldsymbol{y}}\sim p(\bar{\boldsymbol{y}}|\bar{\boldsymbol{z}})}[-\log p(\bar{\boldsymbol{y}}|\bar{\boldsymbol{z}})]+\mathbb{E}_{\bar{\boldsymbol{z}}\sim p(\bar{\boldsymbol{z}})}[-\log p(\bar{\boldsymbol{z}})].$ (12)
We derive the rate loss from this upper bound as:
$R=-\frac{1}{n_{\rm p}n_{\rm c}d_{\rm v}}(\log p(\bar{\boldsymbol{y}}|\bar{\boldsymbol{z}})+\log p(\bar{\boldsymbol{z}})).$ (13)